OpenStudy (iwanttogotostanford):

Please HELP: Using the completing-the-square method, rewrite f(x) = x^2 + 4x − 1 in vertex form. f(x) = (x + 2)^2 + 1 f(x) = (x + 2)^2 f(x) = (x + 2)^2 + 4 f(x) = (x + 2)^2 − 5

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" is? sometimes just called a "perfect square"

OpenStudy (iwanttogotostanford):

yes

OpenStudy (zzr0ck3r):

\[x^2+bx+c=(x+\frac{b}{2})^2+c-\frac{b^2}{4}\]

OpenStudy (iwanttogotostanford):

would it be D then?

OpenStudy (jhannybean):

\[f(x) = x^2+4x-1\]\[=(x^2+4x)-1\]\[=(x^2+4x+\color{red}{4})-1\color{red}{-4}\]\[=\boxed{(x-2)^2-5}\]

OpenStudy (jdoe0001):

so... \(\bf f(x)=x^2+4x-1\implies f(x)=(x^2+4x+{\color{red}{ \square ?}}^2)-1\) what do you think is the missing number?

OpenStudy (jhannybean):

Lovely.

OpenStudy (zzr0ck3r):

correct @iwanttogotostanford

OpenStudy (iwanttogotostanford):

one more?

OpenStudy (jhannybean):

@jdoe0001 why isn't `\boxed` working for me?

OpenStudy (iwanttogotostanford):

Which of the following values "completes the square," or creates a perfect square trinomial, for x2 + 6x + ___? 1 3 6 9

OpenStudy (iwanttogotostanford):

i believe its 9 but i just want to make sure

OpenStudy (jhannybean):

\[c=\left(\frac{6}{2}\right)^2 = ~?\]

OpenStudy (iwanttogotostanford):

9!

OpenStudy (jhannybean):

\[\color{red}{\checkmark}\]

OpenStudy (iwanttogotostanford):

i really need help with this one real quick, thanks for all your help..:-)

OpenStudy (iwanttogotostanford):

Solve x2 + 8x − 3 = 0 using the completing-the-square method. x = four plus or minus the square root of three x = negative four plus or minus the square root of three x = four plus or minus the square root of nineteen x = negative four plus or minus the square root of nineteen

OpenStudy (jhannybean):

use the method we used above.

OpenStudy (iwanttogotostanford):

so , B?

OpenStudy (jhannybean):

I don't know,I didn't solve it.

OpenStudy (iwanttogotostanford):

could you help me? :_)

OpenStudy (jhannybean):

first add +3 to both sides.

OpenStudy (jhannybean):

what do you get, @iwanttogotostanford ?

OpenStudy (iwanttogotostanford):

the same but on the left -6 and on the right 3

OpenStudy (jhannybean):

??....

OpenStudy (jhannybean):

okay, if you add +3 to both sides , the answer i was looking for is, x\(^2\)+8x=3

OpenStudy (iwanttogotostanford):

@Jhannybean sorry, i was gone